OpenStudy (anonymous):

The perimeter of a rectangle is nine times its width. If the width is 20, the length is?

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Actually the book gave the answer as 70 ??

OpenStudy (anonymous):

lol i am starting to think the same thing!

OpenStudy (anonymous):

hey dude this is super wrong... the actual answer is: P=9w P=9(20)= 180 P=2L+2W 180=2L+ 2(20) 2L=140 L=70

OpenStudy (anonymous):

Hey dude, I go to UCI so I got this ;) forget that guy who told you not to listen to your book